# Empirically Exploring How Novices Write Software Models in Alloy

### v2

@@ 1. Introduction
Alloy (JacksonAlloyBook2006, ) is a relational modeling language. A key strength of Alloy is the ability to develop models in the Analyzer, an automatic analysis engine based on off-the-shelf SAT solvers, which the Analyzer uses to generate scenarios that highlight how the modeled properties either hold or are refuted, as desired. The user is able to iterate over these scenarios one by one, inspecting them for correctness. Alloy has been used to verify software system designs (Zave15, ; BagheriETAL2018, ; WickersonETAL2017, ; ChongETAL2018, ), and to perform various forms of analyses over the corresponding implementation, including deep static checking (JacksonVaziri00Bugs, ; TACOGaleottiETALTSE2013, ), systematic testing (MarinovKhurshid01TestEra, ), data structure repair (ZaeemKhurshidECOOP2010, ), automated debugging (GopinathETALTACAS2011, ) and generate security attacks (websecurity, ; Margrave, ; CheckMateMicro2019, ).
However, to gain the many benefits that come from utilizing software models, the model itself needs to be correct. In Alloy, there are two types of faults that can appear in a model: (1) underconstrained faults in which the model allows scenarios it should prevent, and (2) overconstrained faults in which the model prevents scenarios it should allow. To detect an underconstrained fault, during enumeration, the user needs to observe a scenario that they did not expect to see. Unfortunately, even if the user notices something is wrong with a scenario, according to a recent user student, both novice and expert Alloy modelers struggled to refine a faulty formula given an incorrectly produced scenario (alloyformalise23study, ). To detect an overconstrained fault, after enumeration, the user needs to realize that an expected scenario was never present. Since Alloy commands can produce hundreds of scenarios, this places a high burden on the user to realize their Alloy model is incorrect.
@@
Given these difficulties, there is a growing body of work looking at how to detect and debug faulty Alloy models. Initially, a unit testing framework AUnit, was created (AUnitConcept, ; AUnit, ). AUnit enables users to outline a specific scenario they expect their model to allow or prevent and then check that this behavior actually occurs. AUnit has been used for mutation testing (MuAlloyTool, ; MuAlloyTemporal, ), fault localization (AlloyFL, ), and automated repair (ARepair, ; icebar, ). Since then, fault localization and repair of Alloy models has become an active research area (atr, ; flack, ; beafix, ; tar, ). As this field grows, we believe it is important to understand what mistakes new developers actually make when writing models, so that these techniques can (1) be effective for the majority of mistakes made, (2) be evaluated with respect to real faults and not simulated faults and (3) be properly leveraged to ease the burden of learning to write software models. To achieve these goals, this paper presents an empirical study that explores the common development practices of novice Alloy modelers using 97,755 models submitted through Alloy4Fun, an online educational website for Alloy.
@@
Benchmarks for Alloy: As part of our efforts to explore the submissions to Alloy4Fun, we create two benchmarks: (1) a collection of real world faulty models broken down by type of fault and (2) a collection of models that tracks incremental changes made as a user refines a model, which can be used to evaluate and improve incremental analysis techniques for Alloy (Reach, ; iAlloy, ; titanium, ; Platinum, ).
Open Source: We release our benchmarks and organized data analysis at: https://github.com/NoviceAlloyModels
@@ 2. Background
In this section, we describe key concepts of Alloy and Alloy4Fun.
@@ 2.1. Alloy
Figure 1 (a) displays a faulty temporal model of a file system trash can from the Alloy4Fun dataset (alloy4funbenchmark, ). Signature paragraphs introduce named sets and can define relations, which outline relationships between elements of sets. Line 1 introduces a named set File and establishes that each File atom connects to zero or one (lone) File atoms through the link relation. Lines 2 and 3 introduce the named sets Trash and Protected as subsets (in) of File. Signatures and relations can be declared mutable (var), which means that the elements of these sets can vary across different states in the same scenario. In our example, all 3 signatures (File, Trash, and Protected) and the one relation (link) are mutable.
@@ 2.2. Alloy4Fun Exercises
1. // a protected file is at no time sent to the trash 2. pred prop9 {}3. //SECRET4. pred prop9o { always no Protected & Trash }5. check prop9Ok {6.    (prop9 and not prop9o) implies (prop9o iff prop9)7. }
@@ 3. Experiment SetUp
In this paper, we will use the following terminology. A submission consists of the base model, the user’s current attempt for a specific exercise, and any helper predicates invoked. An edit path is the series of submissions a user makes starting with the starter model and ending with their last submission. Edit paths can contain submissions across multiple different exercises but are all within the same model. An attempt is the the series of submissions within an edit path that the user makes for a specific exercise.
@@ 3.1. Experimental Data
In our study, we use the publicly released Alloy4Fun dataset, which contains real world models obtained from student submissions in the academic period from Fall 2019 to Spring 2023 (alloy4funbenchmark, ). In total, there are 97,755 submissions that span 17 different Alloy models and 183 predicates to be filled in (exercises). Of these, we filtered out 2,024 submissions that change the model’s structure or are only changes to the theme settings for visualization. In addition, we filtered 2,448 submissions that are empty. Empty submissions are underconstrained faults. However, since empty submissions do not fundamentally illustrate any effort by the novice user, we removed them to avoid skewing the conclusions. As a result, we end up with 93,283 submissions, all of which have tangible user created content.
Table 1 gives an overview of the complexity of the models used in our study in terms of the universe of discourse each model creates. Column #Sig is the total number of signatures in the model, #Abs is the number of abstract signatures, #Ext is the number of signatures that extend another signature, #Rel is the number of relations, Arity is the average arity of all relations in the model (2 indicates a binary relation), and #Exe is the number of exercises. Models with an underscore in their name represent models that have multiple versions in the dataset. Between versions, the number of exercise, instructional text and/or the type of logic can change.
@@
An exercise falls into one of the following categories: relational logic (RL), predicate logic (PL), first order logic (FOL), and linear temporal logic (LTL). To illustrate the bounds between the different types of logic, the following is the subset of Alloy’s grammar that captures all supported relational logic operators:
@@
The following is the subset of first order logic operators:
@@ 3.2. Classification of Submissions
Within our dataset, a submission to falls into one of the following categories: correct, overconstrained (over), underconstrained (under), both over- and underconstrained (both), type error (type) and syntax error (syntax). Correct submissions are semantically equivalent to the oracle, overconstrained submissions prevent valid scenarios from being generated, underconstrained submissions allow invalid scenarios to be generated, both submissions allow a combination of over- and underconstrained behaviors, type errors are formulas that are syntactically valid but fail to typecheck, and syntax errors are formulas that fail to adhere to the grammar rules.
@@
To determine which category a submission falls under, we treat the oracle solution as the ground truth and consider any submission that is semantically equivalent as correct. If not correct, we then user Alloy to check if the submission is overconstrained, underconstrained or both. If the model fails to compile, we use the generated error message to determine if there was a type or a syntax error.
@@ 3.3. Recreating Submission Revision Histories
When a student checks an exercise, Alloy4Fun logs the student submission with a unique id key and a “derivationOf” parameter that contains the unique id of the parent entry. If the submission is the first attempt, the parent model is the officially published Alloy4Fun example model. Otherwise, the parent is the previous submission made for the exercise. We use this information to re-build edit paths that capture the series of edits a user makes. As an example, the following is the edit path for our example in Figure 1:
@@ 4.1. RQ1: What classification of submissions do novice users make and at what rate?
To get a general overview of the submissions made, Figure 3 shows pie charts depicting the breakdown in classification of submission. The first pie chart contains information for all submissions, while the remaining show the breakdown per type of logic. Submissions were sorted into different types of logic based on the logic in the user’s submission and not the backend oracle solution.
@@
While correct submissions are the largest individual chunk at 25.4%, faulty submissions (both, over and under) combine to account for 45.3% of all submissions. Faults that are both over- and underconstrained account for the largest portion of mistakes at 23.28% of all submissions. However, the distribution of strictly overconstrained or strictly underconstrained faults varies by logic. The rate of underconstrained formulas notably increases for FOL and LTL submissions, going from around 7% for RL and PL to about 15% for FOL and LTL submissions. This indicates that the addition of quantification and temporal logic results in novice users writing properties that are often too permissive. In RQ6, we dive deeper into common mistakes that reveals several reasons why different types of faults are more prevalent in different types of logic.
Invalid submissions combine to account for 29.28% of all submissions and are almost evenly split between syntax errors (14.64%) and type errors (14.65%). While the rate of syntax errors is relatively consistent across types of logic, the rate of type errors varies greatly. Over 15% of RL, PL and FOL submissions have type errors, while only 7% of LTL submissions have type errors. One possible explanation for this is that within the LTL exercises, only some of the signatures and relations are mutable. As a result, there is a more narrow pool of types for which temporal properties apply, resulting in novices making less mistakes.
@@ 4.2. RQ2: How often do novice users make the same submission?
If users often repeat the same submission, then this represents underlying common approaches that novices take to modeling, and understanding what these common approach are can help guide improvements in tool support and educational material. Therefore, we investigate how often new users make duplicate submissions. We consider two forms of equivalence: syntactic and semantic. Syntactical duplicates covey the rate at which novice users literally write the same exact formula, while semantic duplicates allows us to explore the rate at which users express the same underlying concept. To determine if two submissions are syntactically equivalent, we used a PrettyStringVisitor to reprint formulas in a consistent format that eliminates trivial differences in formatting that arise from different uses of white spaces. To determine if two submissions are semantically equivalent, we used the Analyzer to check for logical equivalence between submissions.
Table 3 and Table 4 shows the rate at which submissions are syntactically and semantically unique broken down by classification (Clx) and additionally type of logic for syntactic. For both tables, Column #Sub is the total number of submissions, column #Uni is the number of submissions which are unique syntactically or semantically, %Uni is the percentage of total submissions that are unique. Syntax and Type submissions do not have semantically equivalent values since these submissions do not compile.
In general, novices do have common approaches to modeling – only 55.55% of submissions are syntactically unique and 16.55% are semantically unique. There are two main trends that impact this. First, there is high redundancy in correct submissions, which is expected as there are a finite number of ways to express a property correctly. Moreover, semantically, all correct submissions for an exercise are equivalent. Second, there is notably less redundancy in type and syntax errors, meaning that novice users make a broader range of mistakes when incorrectly structuring a formula compared to writing a syntactically valid formula.
For faulty submissions, RL and PL submissions tend to make the same overly restrictive formula (over) mistakes syntactically, while FOL and LTL submissions tend to repeat the same overly permissive mistakes (under). As RQ6 will further highlight, one reason for this split phenomenon is that when novice users write incorrect relational formulas, they tend to write a formula without quantification, which will apply a concept too broad to an entire set of atoms, overconstraining the model. In contrast, when writing a quantified formula, the user will apply a formula to too narrow of a domain, underconstraining the model. Semantically, users are almost always making the same underconstrained fault in which only 9.83% of the formulas are unique logically. However, with at most 36.17% of the formulas being unique across all classifications of mistakes, users often repeat the same high level logical mistakes.
@@ 4.3. RQ3: How effective is the Analyzer’s compilation error reports?
Nearly a third of the time, a novice user will create a submission that fails to compile. Therefore, we investigate the effectiveness of Alloy’s current compiler-based error reporting, which uses its built in parser plus an internal type system to report syntax and type errors to the user. This effort is summarized in Table 5. For each model, column 2 reports the percentage of attempts that have at least one invalid submission and column 3 presents the percentage of attempts that never turned into valid formulas. The remaining three columns aim to quantify how a user goes from an invalid submission to a valid submission. We first collected all invalid subpaths, which are the portions of attempts that consist of only type or syntax errors edits. Then, column 4 reports the average length, column 5 reports the percentage of subpaths that have a length of 5 or more, and column 6 reports the max length of any subpath. Since an attempt often represents more than one submission and the user can switch between types of logic as they make edits, we do not report information based on type of logic.
@@
Overall, the error reports help novices most of the time. On average, users correct an error within 2 attempts. However, there are gaps in the effectiveness of the current reports: 2.88% of the time, users give up without creating a formula that compile and 7.48% of the time, the user makes 5 or more edits to correct an error. Upon investigating some of these longer invalid subpaths, we found Alloy’s error reporting to be pretty effective for syntax errors but ambigious at best, misleading at worst, for type errors.
Alloy formulas either produce sets or booleans. Often times, a novice’s type mistake is producing a set when a boolean is expected, or vice versa. However, Alloy’s type checking error reports often fail to make this clear. To demonstrate, consider the following chain of edits for a classroom_fol inv15 attempt in which the user starts with a subformula that evaluates to a boolean when a set is expected:
@@
The red highlights show the guidance given by Alloy’s error report. For line 1, the user is only told ‘‘the formula fails to typecheck” with the unary operator some the only highlight. This report fail to inform the user that “Teacher in p.Tutors” evaluates to a boolean and “some” expects a set, leading to the type issue. For lines 2 and 3, the text of the error report – “This must be a set or relation. Instead, it has the following possible type(s): PrimitiveBoolean.” – is more robust. However, this report message is based on the idea that the core formula is supposed to be “a in b,” even though the user could be trying to write “some a.” Moreover, the highlighted formula is actually a valid type checked formula, which introduces confusion. As this chain of edits shows, the user did not pick up on the core of the type issue, as the user only tried to insert parenthesis to change the scope without restructuring the formula.
@@
With each edit, the user is informed that the red highlighted text must “be a formula expression.,” There is no definition of a “formula expression” and no mention of the expectation is that this formula needs to evaluate to a boolean. As a result, this user keeps trying to edit the highlighted formula as if the type error is within the formula itself. The user fails to infer that the quantified formula expects to check whether each atom in the domain is true or false over a boolean subformula not a set subformula.

### v3

# Right or Wrong – Understanding How Novice Users Write Software Models

@@ 1. Introduction
Alloy (JacksonAlloyBook2006, ) is a relational modeling language. A key strength of Alloy is the ability to develop models in the Analyzer, an automatic analysis engine based on off-the-shelf SAT solvers, which the Analyzer uses to generate scenarios that highlight how the modeled properties either hold or are refuted, as desired. The user is able to iterate over these scenarios one by one, inspecting them for correctness. Alloy has been used to verify software system designs (Zave15, ; BagheriETAL2018, ; WickersonETAL2017, ; ChongETAL2018, ), and to perform various forms of analyses over the corresponding implementation, including deep static checking (JacksonVaziri00Bugs, ; TACOGaleottiETALTSE2013, ), systematic testing (MarinovKhurshid01TestEra, ), data structure repair (ZaeemKhurshidECOOP2010, ), automated debugging (GopinathETALTACAS2011, ) and to generate security attacks (websecurity, ; Margrave, ; CheckMateMicro2019, ).
However, to gain the many benefits that come from utilizing software models, the model itself needs to be correct. In Alloy, there are two types of faults that can appear in a model: (1) underconstrained faults in which the model allows scenarios it should prevent, and (2) overconstrained faults in which the model prevents scenarios it should allow. To detect an underconstrained fault, during enumeration, the user needs to observe a scenario that they did not expect to see. Unfortunately, even if the user notices something is wrong with a scenario, according to a recent user study, both novice and expert Alloy modelers struggled to refine a faulty formula given an incorrectly produced scenario (alloyformalise23study, ). To detect an overconstrained fault, after enumeration, the user needs to realize that an expected scenario was never present. Since Alloy commands can produce hundreds of scenarios, this places a high burden on the user to realize their Alloy model is incorrect.
@@
Given these difficulties, there is a growing body of work looking at how to detect and debug faulty Alloy models. Initially, a unit testing framework AUnit, was created (AUnitConcept, ; AUnit, ). AUnit enables users to outline a specific scenario they expect their model to allow or prevent and then check that this behavior actually occurs. AUnit has been used for mutation testing (MuAlloyTool, ; MuAlloyTemporal, ), fault localization (AlloyFL, ), and automated repair (ARepair, ; icebar, ). Since then, fault localization and repair of Alloy models has become an active research area (atr, ; flack, ; beafix, ; tar, ). As this field grows, we believe it is important to understand what mistakes new developers actually make when writing models so that these techniques can (1) be effective for the majority of mistakes made, (2) be evaluated with respect to real faults and not simulated faults and (3) be properly leveraged to ease the burden of learning to write software models. To achieve these goals, this paper presents an empirical study that explores the common development practices of novice Alloy modelers using 97,755 models submitted through Alloy4Fun, an online educational website for Alloy.
@@
Benchmarks for Alloy: As part of exploration, we create two benchmarks: (1) a collection of real-world faulty models broken down by type of fault and (2) a collection of models that tracks changes, which can be used to evaluate and improve incremental analysis techniques for Alloy (Reach, ; iAlloy, ; titanium, ; Platinum, ).
Open Source: We release our benchmarks and data analysis at: https://github.com/NoviceAlloyModels/NoviceAlloyStudy.
@@ 2. Background
In this section, we describe the key concepts of Alloy and Alloy4Fun.
@@ 2.1. Alloy
Figure 1 (a) displays a faulty temporal model of a file system trash can from the Alloy4Fun dataset (alloy4funbenchmark, ). Signature paragraphs introduce named sets and can define relations, which outline relationships between elements of sets. Line 1 introduces a named set File and establishes that each File atom connects to zero or one (lone) File atoms through the link relation. Lines 2 and 3 introduce the named sets Trash and Protected as subsets (in) of File. Signatures and relations can be declared mutable (var), which means that the elements of these sets can vary across different states in the same scenario. In our example, all signatures and relations are mutable.
@@ 2.2. Alloy4Fun Exercises
1. // a protected file is at no time sent to the trash
2. pred prop9 {}
3. //SECRET
4. pred prop9o { always no Protected & Trash }
5. check prop9Ok {
6.    (prop9 and not prop9o) implies (prop9o iff prop9)
7. }
@@ 3. Experiment SetUp
In this paper, we will use the following terminology. A submission consists of the base model, the user’s current attempt for a specific exercise, and any helper predicates invoked. An edit path is the series of submissions a user makes starting with the starter model and ending with their last submission. Edit paths can contain submissions across multiple different exercises but are all within the same model. An attempt is the series of submissions within an edit path that the user makes for a specific exercise.
@@ 3.1. Experimental Data
In our study, we use the publicly released Alloy4Fun dataset, which contains real-world models obtained from student submissions in the academic period from Fall 2019 to Spring 2023 (alloy4funbenchmark, ). In total, there are 97,755 submissions that span 17 different Alloy models and 183 predicates to be filled in (exercises). Of these, we filtered out 2,024 submissions that change the model’s structure or are only changes to the theme settings for visualization. In addition, we filtered 2,448 submissions that are empty. Empty submissions are underconstrained faults. However, since empty submissions do not fundamentally illustrate any effort by the novice user, we removed them to avoid skewing the conclusions. As a result, we end up with 93,283 submissions, all of which have tangible user-created content.
Table 1 gives an overview of the complexity of the models used in our study in terms of the universe of discourse each model creates. Column #Sig is the total number of signatures in the model, #Abs is the number of abstract signatures, #Ext is the number of signatures that extend another signature, #Rel is the number of relations, Arity is the average arity of all relations in the model (2 indicates a binary relation), and #Exe is the number of exercises. Models with an underscore in their name represent models that have multiple versions in the dataset. Between versions, the number of exercises, instructional text, and/or the type of logic can change.
@@
An exercise falls into one of the following categories: relational logic (RL), predicate logic (PL), first-order logic (FOL), and linear temporal logic (LTL). The following is the subset of Alloy’s grammar that captures all supported relational logic operators:
@@
The following is the subset of first-order logic operators:
@@ 3.2. Classification of Submissions
Within our dataset, a submission falls into one of the following categories: correct, overconstrained (over), underconstrained (under), both over- and underconstrained (both), type error (type), and syntax error (syntax). Correct submissions are semantically equivalent to the oracle, overconstrained submissions prevent valid scenarios from being generated, underconstrained submissions allow invalid scenarios to be generated, both submissions allow a combination of over- and underconstrained behaviors, type errors are formulas that are syntactically valid but fail to typecheck, and syntax errors are formulas that fail to adhere to the grammar rules.
@@ 3.3. Recreating Submission Revision Histories
When a student checks an exercise, Alloy4Fun logs the student submission with a unique id key and a “derivationOf” parameter that contains the unique id of the parent entry. If the submission is the first attempt, the parent model is the officially published Alloy4Fun example model. Otherwise, the parent is the previous submission made by the user. We use this information to re-build edit paths that capture the series of edits a user makes. As an example, the following is the edit path for our example in Figure 1:
@@
For edit paths, we do include empty submissions. While empty submissions do not hold much value as individual submissions, in the middle of an edit path, an empty submission can convey when users reset their attempt on an exercise.
@@ 4.1. RQ1: What classification of submissions do novice users make and at what rate?
To get a general overview of the submissions made, Figure 3 shows pie charts depicting the breakdown in the classification of submissions. The first pie chart contains information for all submissions, while the remaining shows the breakdown per type of logic. Submissions were sorted into different types of logic based on the logic in the user’s submission and not the backend oracle solution.
@@
While correct submissions are the largest individual chunk at 25.4%, faulty submissions (both, over and under) combine to account for 45.3% of all submissions. Faults that are both over- and underconstrained account for the largest portion of mistakes at 23.28% of all submissions. However, the distribution of strictly overconstrained or strictly underconstrained faults varies by logic. The rate of underconstrained formulas notably increases for FOL and LTL submissions, going from around 7% for RL and PL to about 15% for FOL and LTL submissions. This indicates that the addition of quantification and temporal logic results in novice users writing properties that are often too permissive. In latter questions, we dive deeper into why different types of faults occur.
Invalid submissions combine to account for 29.28% of all submissions and are almost evenly split between syntax errors (14.64%) and type errors (14.65%). While the rate of syntax errors is relatively consistent across types of logic, the rate of type errors varies greatly. Over 15% of RL, PL, and FOL submissions have type errors, while only 7% of LTL submissions have type errors. One possible explanation for this is that within the LTL exercises, only some of the signatures and relations are mutable. As a result, there is a more narrow pool of types for which temporal properties apply, resulting in novices making fewer mistakes.
@@ 4.2. RQ2: How often do novice users make the same submission?
If users often repeat the same submission, then this represents underlying common approaches that novices take to modeling, and understanding what these common approaches are can help guide improvements in tool support and educational material. Therefore, we investigate how often new users make duplicate submissions. We consider two forms of equivalence: syntactic and semantic. To determine if two submissions are syntactically equivalent, we used a PrettyStringVisitor to reprint formulas in a consistent format that eliminates trivial differences in formatting that arise from different uses of white spaces. To determine if two submissions are semantically equivalent, we used the Analyzer to check for logical equivalence between submissions.
Table 3 and Table 4 show the rate at which submissions are syntactically and semantically unique broken down by classification (Clx) and additionally type of logic for syntactic. For both tables, Column #Sub is the total number of submissions, column #Uni is the number of submissions that are unique syntactically or semantically, %Uni is the percentage of total submissions that are unique. Syntax and Type submissions do not have semantically equivalent values since these submissions do not compile.
In general, novices do have common approaches to modeling – only 55.55% of submissions are syntactically unique and 16.55% are semantically unique. Two main trends impact this. First, there is high redundancy in correct submissions, which is expected as there are a finite number of ways to express a property correctly. Moreover, semantically, all correct submissions for an exercise are equivalent. Second, there is notably less redundancy in type and syntax errors, meaning that novice users make a broader range of mistakes when incorrectly structuring a formula compared to writing a syntactically valid formula that is wrong.
For faulty submissions, RL and PL submissions tend to make the same overly restrictive formula (over) mistakes syntactically, while FOL and LTL submissions tend to repeat the same overly permissive mistakes (under). As RQ6 will further highlight, one reason for this split phenomenon is that when novice users write incorrect relational formulas, they tend to write a formula that checks for a concept too broadly across an entire set of atoms when it should be checked per atom of a set, overconstraining the model. In contrast, when writing a quantified formula, the user will apply a formula to too narrow of a domain, underconstraining the model. Semantically, users are almost always making the same underconstrained fault, as only 9.83% of the formulas are unique logically. However, with at most 36.17% of the formulas being unique across all classifications of mistakes, users often repeat the same high-level logical mistakes.
@@ 4.3. RQ3: How effective is the Analyzer’s compilation error reports?
Nearly a third of the time, a novice user will create a submission that fails to compile. Therefore, we investigate the effectiveness of Alloy’s current compiler-based error reporting, which uses its built-in parser plus an internal type system to report syntax and type errors to the user. This effort is summarized in Table 5. For each model, column 2 reports the percentage of attempts that have at least one invalid submission, and column 3 presents the percentage of attempts that never turned into valid formulas. The remaining three columns aim to quantify how a user goes from an invalid submission to a valid submission. We first collected all invalid subpaths, which are the portions of attempts that consist of only type or syntax error edits. Then, column 4 reports the average length, column 5 reports the percentage of subpaths that have a length of 5 or more, and column 6 reports the max length of any subpath. Since an attempt often represents more than one submission and the user can switch between types of logic as they make edits, we do not report information based on the type of logic.
@@
Overall, the error reports help novices most of the time. On average, users correct an error within 2 attempts. However, there are gaps in the effectiveness of the current reports: 2.88% of the time, users give up without creating a formula that compiles and 7.48% of the time, the user makes 5 or more edits to correct an error. Upon investigating some of these longer invalid subpaths, we found Alloy’s error reporting to be pretty effective for syntax errors but ambiguous at best, and misleading at worst, for type errors.
Alloy formulas either produce sets or booleans. Oftentimes, a novice’s type mistake is producing a set when a boolean is expected, or vice versa. However, Alloy’s type-checking error reports often fail to make this clear. To demonstrate, consider the following chain of edits for a classroom_fol inv15 attempt in which the user starts with a subformula that evaluates to a boolean when a set is expected:
@@
The highlights show the guidance given by Alloy’s error report. For line 1, the user is only told “the formula fails to typecheck” with the unary operator some the only highlight. This report fails to inform the user that “Teacher in p.Tutors” evaluates to a boolean and “some” expects a set, leading to the type issue. For lines 2 and 3, the text of the error report – “This must be a set or relation. Instead, it has the following possible type(s): PrimitiveBoolean.” – is more robust. However, the highlighted formula “some Teacher” is actually a valid type checked formula, which introduces confusion. As this chain of edits shows, the user did not pick up on the core of the type issue, as the user kept trying to solve the problem by inserting parenthesis but only changed the type error from “some [boolean]” to “[boolean] in [set].”
@@
With each edit, the user is informed that the highlighted text must “be a formula expression.,” There is no definition of a “formula expression” and no mention of the expectation that this formula needs to evaluate to a boolean. As a result, this user keeps trying to edit the highlighted formula as if the type error is within the formula. However, the user fails to realize that the edited formulas typecheck individually, but the quantified formula encompassing these formulas expects a boolean subformula not a set subformula.